# The equilibrium properties of obvious strategy profiles

### v1

@@ Abstract
This paper studies the equilibrium properties of the “obvious strategy profile” in large finite-player games. Each player in such a strategy profile simply adopts a randomized strategy as she would have used in a symmetric equilibrium of an idealized large game. We show that, under a continuity assumption, (i) obvious strategy profiles constitute a convergent sequence of approximate symmetric equilibria as the number of players tends to infinity, and (ii) realizations of such strategy profiles also form a convergent sequence of (pure strategy) approximate equilibria with probability approaching one. Our findings offer a solution that is easily implemented without coordination issues and is asymptotically optimal for players in large finite games. Additionally, we present a convergence result for approximate symmetric equilibria.
@@
Keywords: Large games, large finite games, auxiliary mapping, obvious strategy profile, symmetric equilibrium, approximate equilibrium.
@@ 1 Introduction
Games with a continuum of agents (referred to as large games) have been extensively studied in the literature. Typically, such games model strategic interactions involving a large but finite number of participants. By assuming a continuum of players where each player’s unilateral deviation does not affect the aggregate action distribution, various desirable properties, such as the existence of pure strategy Nash equilibrium, have been established in these idealized large games. Despite significant progress in recent decades, the predominant focus in existing literature has been on the existence and convergence of pure strategy Nash equilibria in large games. However, an important question remains unanswered: do players truly benefit from such an idealized process? Specifically, while players may acknowledge that the Nash equilibrium of the idealized large game predicts a desirable social choice outcome, how do they behave to implement such an outcome in their real-life (finite) games? Furthermore, given an implementation strategy profile, do players have incentives to adhere to this designated strategy profile?
@@
The first question seems to have a straightforward answer: each player could simply adopt the strategy they would select in an idealized large game. For instance, if a Nash equilibrium in the large game requires every player to choose a common strategy, then players in the finite real game can follow that equilibrium strategy. While this approach works in some situations, it often fails due to coordination issues arising from multiple individual-optimal choices.
To illustrate this, consider a simple example involving many players in a holiday village who must decide between going to the beach (action $B$) or staying at the hotel (action $H$). For each player, the payoff for staying at the hotel is always 00. Conversely, if a player chooses to go to the beach, her payoff depends on the proportion of players who also choose the beach. Specifically, her payoff is positive if only a few players go to the beach; however, as more players join her there, her payoff decreases and eventually turns negative once the number of beachgoers becomes sufficiently large. Formally, let $I$ denote the set of players and $A=\{B,H\}$ represent the action set available to them. Let $\tau(B)$ (resp. $\tau(H)$) denote the proportion of players selecting action $B$ (resp. $H$). Players in this game share a common payoff function given by:
$u(H,\tau)=0,\,\,\,\,\,u(B,\tau)=\tfrac{1}{2}-\tau(B),$
where $\tau=(\tau(B),\tau(H))$ is an action distribution.
This is a large finite-player game, and to simplify the analysis, players assume they are participating in a game with a continuum of players. In this imagined large game, there exists a pure strategy Nash equilibrium where each action is chosen by exactly half of the players, as each player in this scenario is indifferent between the actions $H$ and $B$ [^1] It can be verified that this is the unique pure strategy Nash equilibrium. Therefore, the action distribution $\tau^{*}=(\tfrac{1}{2},\tfrac{1}{2})$ represents the desirable social choice outcome predicted by the large game model. However, issues arise when considering the implementation of this outcome in a real finite game. Although players are aware of the equilibrium action distribution, each player has multiple optimal choices available given this distribution, [^2] Given the equilibrium action distribution $\tau^{*}=(\tfrac{1}{2},\tfrac{1}{2})$, it follows that $u(H,\tau^{*})=u(B,\tau^{*})$, indicating that both $B$ and $H$ are best responses for each player. but they cannot arbitrarily select one action, as doing so would lead to a non-equilibrium outcome. Thus, if players in the real finite game aim to use pure strategies to achieve the outcome $(\tfrac{1}{2},\tfrac{1}{2})$, they must either communicate with one another before the game begins or rely on a social planner to divide them into two groups and designate an action for each group. However, both of these approaches are unrealistic in scenarios involving many players.
@@
We address this coordination issue in two steps. Firstly, we observe that players in a game are distinguished by their characteristics, typically represented by their payoff functions. For example, drivers in a large traffic network are differentiated by the types of vehicles they drive, each associated with distinct travel times (i.e., different payoff functions). Similarly, buyers in a large online market can be categorized into groups based on their preferences or payoffs for different commodities. Thus, the coordination issue can be resolved if players make decisions based on their characteristics, allowing those with the same characteristic to make the same choice. Such an idea is captured by the notion of symmetry. Therefore, the first step of our approach is to symmetrize the equilibrium outcome. Specifically, given an equilibrium action distribution $\tau^{*}$, we can transform it into a symmetric Nash equilibrium $f^{*}$ using an auxiliary mapping. This mapping ensures that (i) players sharing the same characteristic (i.e., the same payoff function) choose the same strategy in $f^{*}$, and (ii) the action distribution of $f^{*}$ equals $\tau^{*}$.
Secondly, we introduce the concept of an “obvious strategy profile” in real finite-player games. In this strategy profile, players make decisions based on their characteristics, where each player selects a strategy associated with her characteristic in the symmetric equilibrium $f^{*}$ of the imagined large game. For illustration, consider the Beach vs. Hotel game introduced earlier. Since all players in this game share a common payoff function, they possess the same characteristic and consequently choose the same strategy in the equilibrium $f^{*}$. Specifically, all players in $f^{*}$ could adopt a simple randomized strategy $\mu^{*}$, choosing each action with equal probability. The law of large numbers then ensures that the realized action distribution from this symmetric strategy profile matches the equilibrium action distribution, with each action chosen by half of the players. Finally, in real-life finite scenarios, the obvious strategy profile assigns $\mu^{*}$ to each player.
@@
Given the implementation via symmetrization and obvious strategy profile described above, a natural question arises: do players in the real-life finite game have an incentive to follow the obvious strategy profile? This question pertains to the equilibrium property of such a strategy profile. In our motivating example above, the answer is clear: obvious strategy profile forms an approximate Nash equilibrium: suppose that the number of players is $n$. Based on some combinatorial calculations, we can determine that each player’s expected payoff when choosing action $B$ (resp. action $H$) is $-\frac{1}{2n}$ (resp. 00). [^3] Since there are $n$ players in this game and each of them takes a proportion of $\frac{1}{n}$, the expected payoff for each player is $\sum_{k=0}^{n-1}(\tfrac{1}{2}-\tfrac{k+1}{n})\binom{n-1}{k}(\tfrac{1}{2})^{n-1}=\tfrac{1}{2}-\frac{n+1}{2n}=-\tfrac{1}{2n}$ by choosing action $B$ , and 00 by choosing action $H$. Therefore, the obvious strategy profile adopted by all the players is a $\tfrac{1}{4n}$-Nash equilibrium in the $n$-player game. [^4] In an $\varepsilon$-Nash equilibrium, a large portion of players (more than $1-\varepsilon$) choose strategies that are within $\varepsilon$ of their optimal payoffs; see Definition 4 below. Moreover, as $n$ tends to infinity, we observe a convergent sequence of approximate equilibria, indicating that adopting the obvious strategy profile is asymptotically optimal for each player.
@@
However, the asymptotic optimality of obvious strategy profile does not hold in general situations. Example 1 in Section 3 presents a large game $G$ with a Nash equilibrium $g$; however, the obvious strategy profiles induced from $g$ do not form a convergent sequence. As the main result of this paper, we reestablish this convergence result by focusing on a class of large games that satisfy a continuity condition. In particular, Theorem 1 shows that for any large game with a convergent sequence of finite-player games, under this continuity condition, the sequence of obvious strategy profiles forms a convergent sequence of approximate symmetric equilibria. Moreover, such a continuity condition is satisfied in most large games discussed in the literature, including large games with finite characteristics, and continuous large games with continuous equilibria.
@@
Since an obvious strategy profile typically involves randomized strategies, Theorem 1 only guarantees the ex ante asymptotic optimality of such profiles. Therefore, it is natural to consider the ex post equilibrium property of the obvious strategy profile after the resolution of uncertainty. As the second main result of this paper, Theorem 2 shows that, under the same continuity condition, the sequence of realized obvious strategy profiles forms a convergent sequence of approximate equilibria, with probability approaching one. Therefore, based on the concept of obvious strategy profile and the main theorems, we have provided a complete answer to our initial question: players in a large finite game benefit from the idealization process by playing the obvious strategy profile.
@@
Our main results provide a theoretical explanation for human behavior in large strategic interactions. For example, people increasingly use traffic apps, such as Google Maps, to select routes before traveling. In most traffic apps, a driver must input an origin point, a destination point, and the type of vehicle they drive (e.g., car, bus, or truck). The app then calculates the expected travel times for different feasible paths based on idealized models and historical traffic data. Finally, it recommends several optimal choices to the driver based on its calculations and the vehicle type. Typically, a driver will randomly select among the optimal choices recommended by the app, which can be considered an obvious strategy profile. As there are many drivers in the traffic network, each driver’s choice becomes asymptotically optimal not only ex ante but also ex post, with high probability. [^5] There are many other related examples beyond path selection, such as buyers’ behavior in a large online marketplace. For additional examples of symmetric equilibrium in large economies, see Tirole (1988) and Spiegler (2006, 2016).
Then we introduce some results related to Theorem 1. Proposition 1 demonstrates that for a sequence of finite-player games converging to a large game $G$, every Nash equilibrium of $G$ can be approximated by a sequence of approximate symmetric equilibria from the corresponding finite games. Moreover, this result cannot be strengthened by replacing approximate equilibria with exact equilibria. We also examine obvious strategy profiles in pure strategies; Corollary 2 shows that under specific conditions, these obvious strategy profiles exist in pure strategies and form a convergent sequence.
@@
The remainder of this paper is organized as follows. In Section 2, we introduce models of large games and finite-player games, along with the concepts of Nash equilibrium and approximate equilibrium. Our main results are presented in Section 3. Section 4 discuss some related results and reviews related literature. Finally, all the proofs are collected in Section 5.
@@ 2 The basic model
In this section, we introduce some notations and basic definitions of large games and large finite-player games. In such games, all the players have a common compact set of actions, and each player’s payoff depends on her own choice as well as the action distribution induced by all the players’ choices.
@@ 2.1 Large games
A large game is defined as follows. Let $(I,\mathcal{I},\lambda)$ be an atomless probability space denoting the set of players, [^6] Throughout this paper, we follow the convention that a probability space is complete and countably additive. and $A$ be a compact metric space representing a common action space endowed with the Borel $\sigma$-algebra $\mathcal{B}(A)$. [^7] To simplify the analysis, we restrict attention to large games with common action spaces. All the main results in this paper can be generalized to encompass the setting where players have different feasible action sets. See Remark 1 for more details. The set of Borel probability measures on $A$ is denoted by $\mathcal{M}(A)$. Given an action profile of all the players, the action distribution that specifies the portions of players taking some actions in $A$ (also called a societal summary) can be viewed as an element in $\mathcal{M}(A)$. Each player’s payoff is a bounded continuous function on $A\times\mathcal{M}(A)$, which means that the payoff function continuously depends on her own choice and societal summary.
Let $\mathcal{U}_{A}$ be the space of bounded continuous functions on $A\times\mathcal{M}(A)$ endowed with the sup-norm topology and the resulting Borel $\sigma$-algebra. In a general large game, a player’s characteristic comprises her feasible action set and her payoff function $u_{i}$ (an element of $\mathcal{U}_{A}$). Since players have a common action set $A$ in our setting, $\mathcal{U}_{A}$ can be viewed as the characteristic space for all the players.
A large game $G$ is a measurable function from $(I,\mathcal{I},\lambda)$ to $\mathcal{U}_{A}$, which assigns a payoff function to each player. A pure strategy profile $f$ is a measurable function from $(I,\mathcal{I},\lambda)$ to $A$. Let $\lambda f^{-1}$ be the societal summary (also denoted by $s(f)$), which is the societal action distribution induced by $f$. That is, $s(f)(B)$ ($B$ is a subset of $A$) is the portion of players taking actions in $B$.
[…]
We can now present the equilibrium concepts in large games. The formal definition of a randomized strategy Nash equilibrium is as follows.
@@
A randomized strategy profile $g\colon I\rightarrow\mathcal{M}(A)$ is said to be a randomized strategy Nash equilibrium if for $\lambda$-almost all $i\in I$,
[…]
Therefore, a randomized strategy profile $g$ is a randomized strategy Nash equilibrium if it is optimal for almost every player with respect to the societal summary $s(g)$ in terms of expected payoff. Since for a pure strategy profile $f$, its average action distribution $s(f)$ reduces to the societal action distribution $\lambda f^{-1}$. We have the following definition of pure strategy Nash equilibrium.
@@
A pure strategy profile $f\colon I\rightarrow A$ is said to be a pure strategy Nash equilibrium if, for $\lambda$-almost all $i\in I,$
@@
In the definitions of randomized and pure strategy Nash equilibrium, we only require that almost all players make optimal choices. However, stronger concepts can be established by adjusting the strategies of players with measure zero. While a randomized strategy Nash equilibrium is guaranteed to exist in large games, a pure strategy Nash equilibrium may not exist; see, for example, Qiao and Yu (2014).

### v2

# The equilibrium properties of direct strategy profiles

@@ Abstract
This paper studies the equilibrium properties of the direct strategy profile in large finite-player games. Each player in such a strategy profile simply adopts a strategy as she would have used in a symmetric equilibrium of an idealized large game. We show that, under a mild continuity condition, (i) direct strategy profiles constitute a convergent sequence of approximate equilibria as the number of players tends to infinity, and (ii) realizations of such strategy profiles also form a convergent sequence of (pure strategy) approximate equilibria with probability approaching one. Our findings provide a simple and decentralized approach for implementing equilibrium in large games, yielding outcomes that are asymptotically optimal both ex ante and ex post.
@@
Keywords: Large games, large finite-player games, auxiliary mapping, direct strategy profile, symmetric equilibrium, approximate equilibrium.
@@ 1 Introduction
Games with a continuum of agents (referred to as large games) have been extensively studied in the literature. Typically, such games model strategic interactions involving a large but finite number of participants. By assuming a continuum of players where each player’s unilateral deviation does not affect the aggregate action distribution, various desirable properties, such as the existence of pure strategy Nash equilibrium, have been established in these idealized large games. Despite significant progress in recent decades, the predominant focus in literature has been on the existence and convergence of Nash equilibria in large games. However, an important question remains unanswered: do players truly benefit from such an idealized process? Specifically, while players may acknowledge that the Nash equilibrium of the idealized large game predicts a desirable social choice outcome, how do they behave to implement such an outcome in their real-life (finite) games? Furthermore, given an implementation strategy profile, do players have incentives to adhere to this designated strategy profile?
@@
The first question seems to have a straightforward answer: each player could simply adopt the strategy they would select in an idealized large game. For instance, if a Nash equilibrium in the large game requires every player to choose a common strategy, then players in the real finite game can follow that equilibrium strategy. While this approach works in some situations, it often fails due to coordination issues arising from multiple individual-optimal choices.
To illustrate this, we consider a simple example of routing games. Imagine that many drivers traveling from an origin node $o$ to a terminal node $t$ via two paths, denoted as $a$ and $b$ respectively. The travel time on each path $p$ depends solely on the proportion of drivers using that path, denoted as $\tau(p).$ Path $b$ is a wider route where the travel time matches the proportion of drivers using it. In contrast, Path $a$ is narrower, causing congestion to build up more quickly, resulting in a travel time that is twice the proportion of drivers choosing it, i.e. $2\tau(a)$.
oottPath aaTravel time is $2\tau(a)$Path bbTravel time is $\tau(b)$
This is a large but finite game, and to simplify the analysis, it is natural to consider its continuum-player counterpart. In the corresponding continuum game, there exists a unique Nash equilibrium action distribution in which one-third of the drivers choose the narrower Path $a$, and the remaining two-thirds choose Path $b$. In this equilibrium, both paths yield the same travel time of $\tfrac{2}{3}$, so each driver is indifferent between choosing $a$ and $b$. The resulting action distribution $\tau^{*}=(\tfrac{1}{3},\tfrac{2}{3})$ represents the socially desirable outcome predicted by the continuum model. The challenge lies in implementing $\tau^{*}$ in the actual finite-player setting. Although $\tau^{*}$ specifies the target aggregate behavior, it does not uniquely pin down individual actions: both $a$ and $b$ are best responses. Without coordination, independently chosen actions may fail to reproduce $\tau^{*}$, leading to deviations from equilibrium. Achieving $\tau^{*}$ in pure strategies thus requires communication or centralized assignment—both of which become impractical in large, decentralized environments.
@@
We address this issue in two steps. First, we observe that players in large games are typically distinguished by their characteristics, most notably by their payoff functions. For instance, drivers in a large traffic network may differ by vehicle type, with each type associated with distinct travel times. Similarly, buyers in an online marketplace can be grouped by their preferences over goods. In such environments, coordination becomes feasible if players base their strategies on observable characteristics, so that those with identical characteristics adopt the same strategy. This idea is formalized through the notion of symmetry. The first step of our approach is therefore to construct a symmetric equilibrium. Given an aggregate action distribution $\tau^{*}$ from a Nash equilibrium, we show that it can be implemented by a symmetric strategy profile $f^{*}$ satisfying two properties: (i) players with the same characteristic (i.e., the same payoff function) choose the same strategy in $f^{*}$, and (ii) the aggregate action distribution induced by $f^{*}$ equals $\tau^{*}$.
Second, we introduce the notion of a “direct strategy profile” for the finite-player game. In this strategy profile, each player selects a strategy based on her characteristic by adopting the strategy prescribed for that characteristic in the symmetric equilibrium $f^{*}$ of the continuum model. To illustrate, consider the routing game discussed earlier. Since all drivers share the same payoff function, they are identical in terms of characteristics and therefore adopt the same strategy in the symmetric equilibrium $f^{*}$. For instance, $f^{*}$ may prescribe a simple randomized strategy $\mu^{*}$ that assigns probability $\tfrac{1}{3}$ to Path $a$ and $\tfrac{2}{3}$ to Path $b$. In the corresponding direct strategy profile for the finite game, each player is assigned this strategy $\mu^{*}$.
@@
Given the implementation via symmetrization and the direct strategy profile described above, a natural question arises: do players in the finite game have incentives to follow the direct strategy profile? This question concerns the equilibrium properties of such strategy profiles. In the routing game introduced earlier, the answer is affirmative: the direct strategy profile constitutes an approximate Nash equilibrium. Suppose there are $n$ drivers. Based on a combinatorial argument, [^1] Each driver chooses Path $a$ with probability $\tfrac{1}{3}$ and faces a random number of other drivers making the same choice. The expected travel time for choosing Path $a$ is $2\sum_{k=0}^{n-1}\left(\tfrac{1}{n}+\tfrac{k}{n}\right)\binom{n-1}{k}(\tfrac{1}{3})^{k}(\tfrac{2}{3})^{n-1-k}=\tfrac{2n+4}{3n}$, while the expected travel time for Path $b$ is $\sum_{k=0}^{n-1}\left(\tfrac{1}{n}+\tfrac{n-1-k}{n}\right)\binom{n-1}{k}(\tfrac{1}{3})^{k}(\tfrac{2}{3})^{n-1-k}=\tfrac{2n+1}{3n}$. the expected travel time for a driver choosing Path $a$ is $\tfrac{2n+4}{3n}$, while choosing Path $b$ yields $\tfrac{2n+1}{3n}$. Thus, the direct strategy profile forms a $\tfrac{1}{3n}$-Nash equilibrium [^2] In an $\varepsilon$-Nash equilibrium, a large portion of players (more than $1-\varepsilon$) choose strategies that are within $\varepsilon$ of their optimal payoffs; see Definition 4 below. in the $n$-player game. Moreover, as $n\to\infty$, the approximation error vanishes. This implies that following the direct strategy profile becomes asymptotically optimal in this routing game.
@@
However, the asymptotic optimality of the direct strategy profile does not hold in general. Example 1 in Subsection 4.1 presents a large game $G$ with a Nash equilibrium $g$, for which the associated sequence of direct strategy profiles fails to converge. As the main result of this paper, we recover convergence by restricting attention to a class of large games that satisfy a continuity condition. In particular, Theorem 1 shows that for any large game admitting a convergent sequence of finite-player approximations, the corresponding sequence of direct strategy profiles forms a convergent sequence of approximate symmetric equilibria, provided a continuity condition is satisfied. This condition is relatively mild and holds in many large games studied in the literature, including those with finitely many characteristics and continuous large games with continuous equilibria.
@@
Since a direct strategy profile typically involves randomized strategies, Theorem 1 establishes only ex ante asymptotic optimality. It is therefore natural to ask whether the direct strategy profile remains approximately optimal ex post, after the resolution of uncertainty. As the second main result of this paper, Theorem 2 shows that, under the same continuity condition, the sequence of realized direct strategy profiles forms a convergent sequence of (pure strategy) approximate equilibria, with probability approaching one. Together, the concept of the direct strategy profile and our two main theorems provide a complete answer to the motivating question: players in a large finite game benefit from the idealization process by playing the direct strategy profile.
@@
Our main results offer a theoretical explanation for individual behavior in large strategic environments. For example, people increasingly use traffic apps, such as Google Maps, to select routes before traveling. In most traffic apps, a driver must input an origin point, a destination point, and the type of vehicle they drive (e.g., car, bus, or truck). The app then calculates the expected travel times for different feasible paths based on idealized models and historical traffic data. Finally, it recommends several optimal choices to the driver based on its calculations and the vehicle type. Typically, a driver will randomly select among the optimal choices recommended by the app. This behavior is consistent with the notion of a direct strategy profile. Given the large number of drivers, such decentralized choices become approximately optimal—both ex ante and, with high probability, ex post. Similar reasoning applies to other settings, such as buyer decisions in large online marketplaces. [^3] For further examples of symmetric equilibrium in large economies, see Spiegler (2006), Spiegler (2016), and Tirole (1988).
@@
The remainder of the paper is organized as follows. Section 2 introduces the models of large games and large finite-player games, along with the relevant equilibrium concepts. Our main results are presented in Section 3. Section 4 discusses related results and reviews the relevant literature. Finally, all proofs are collected in Section 5.
@@ 2 The basic model
In this section, we introduce some notation and basic definitions related to large games and large finite-player games. In such games, all players share a common compact action set, and each player’s payoff depends on her own chosen action as well as the action distribution induced by all players’ choices. The definitions in our model follow standard conventions in the literature; see, for example, Khan and Sun (2002).
@@ 2.1 Large games
A large game is defined as follows. Let $(I,\mathcal{I},\lambda)$ be an atomless probability space representing the set of players. [^4] Throughout this paper, we assume that all probability spaces are complete and countably additive. Let $A$ be a compact metric space representing a common action set, equipped with the Borel $\sigma$-algebra $\mathcal{B}(A)$. [^5] To simplify the analysis, we focus on large games with a common action space. All the main results in this paper can be generalized to settings where players have different feasible action sets. See Remark 1 for more details. The set of probability measures on $A$ is denoted by $\mathcal{M}(A)$. Given an action profile for all players, the induced action distribution—which specifies the proportion of players choosing each action in $A$ (also referred to as the societal summary)—can be identified as an element of $\mathcal{M}(A)$. Each player’s payoff is a bounded, continuous function on $A\times\mathcal{M}(A)$, reflecting continuous dependence on both her own action and the societal summary.
Let $\mathcal{U}_{A}$ be the space of bounded continuous functions on $A\times\mathcal{M}(A)$, endowed with the sup-norm topology and the corresponding Borel $\sigma$-algebra. In a general large game, a player’s characteristic consists of her feasible action set and her payoff function $u_{i}$ (an element of $\mathcal{U}_{A}$). Since players share a common action set $A$ in our setting, $\mathcal{U}_{A}$ can be regarded as the characteristic space for all players.
A large game $G$ is a measurable mapping from $(I,\mathcal{I},\lambda)$ to $\mathcal{U}_{A}$, assigning a payoff function to each player. A pure strategy profile $f$ is a measurable function from $(I,\mathcal{I},\lambda)$ to $A$. Let $\lambda f^{-1}$ denote the societal summary (also written as $s(f)$), which is the action distribution induced by $f$. Specifically, for any measurable subset $B\subseteq A$, $[s(f)](B)$ represents the proportion of players choosing actions in $B$.
A randomized strategy for each player $i$ is a probability measure $\mu\in\mathcal{M}(A)$. A randomized strategy profile can be viewed as a measurable mapping $g\mathrel{\mathop{\ordinarycolon}}I\to\mathcal{M}(A)$. [^6] A measurable mapping $g\mathrel{\mathop{\ordinarycolon}}I\to\mathcal{M}(A)$ can also be viewed as a transition probability $g\mathrel{\mathop{\ordinarycolon}}I\times\mathcal{B}(A)\to[0,1]$ such that: (i) For every $B\in\mathcal{B}(A)$, $g(\cdot;B)$ is measurable; (ii) For $\lambda$-almost all $i\in I$, $g(i;\cdot)\in\mathcal{M}(A)$. Note that every pure strategy profile $f$ naturally induces a randomized strategy profile $g^{f}$ defined by $g^{f}(i)=\delta_{f(i)}$ for each player $i\in I$. [^7] Here $\delta_{f(i)}$ denotes the Dirac probability measure that assigns probability one to $\{f(i)\}$. Given a randomized strategy profile $g$, the societal summary $s(g)$ is modeled as the average action distribution across all players, that is, $s(g)=\int_{I}g(i)\,\mathrm{d}\lambda(i)\in\mathcal{M}(A)$. When $A$ is a finite set of actions, $s(g)$ can be identified with a point in the simplex in $\mathbb{R}^{|A|}$. In the case of an infinite action set, $s(g)$ satisfies $s(g)](B)=\int_{I}g(i,B)\,\mathrm{d}\lambda(i)$ for every measurable subset $B\subseteq A$. Moreover, a randomized strategy profile $g$ is said to be symmetric if for any two players $i$ and $i^{\prime}$, we have $g(i)=g(i^{\prime})$ whenever $G(i)=G(i^{\prime})$, i.e., players with the same characteristic (payoff function) use the same strategy.
We now present the equilibrium concepts for large games. The formal definition of a randomized strategy Nash equilibrium is as follows.
@@
A randomized strategy profile $g\colon I\rightarrow\mathcal{M}(A)$ is said to be a randomized strategy Nash equilibrium if for almost all $i\in I$,
$\int_{A}u_{i}\bigl{(}a,s(g)\bigr{)}g(i,\operatorname{d\!}a)\geq\int_{A}u_{i}\bigl{(}a,s(g)\bigr{)}\operatorname{d\!}\mu(a)\text{ for all }\mu\in\mathcal{M}(A).$
Therefore, a randomized strategy profile $g$ is a randomized strategy Nash equilibrium if it is optimal for almost every player with respect to the societal summary $s(g)$, in terms of expected payoff. In the case of a pure strategy profile $f$, the societal summary $s(f)$ coincides with the action distribution $\lambda f^{-1}$. This leads to the following definition of a pure strategy Nash equilibrium.
@@
A pure strategy profile $f\colon I\rightarrow A$ is said to be a pure strategy Nash equilibrium if, for all $i\in I,$
@@
A randomized strategy Nash equilibrium always exists in large games. By contrast, the existence of a pure strategy Nash equilibrium generally requires additional conditions: it is guaranteed in large games with finite action sets (Schmeidler (1973)) or when the player space satisfies certain saturation condition (He et al. (2017)). [^8] See also Qiao and Yu (2014) for counterexamples and further discussion.